OpenStudy (anonymous):

how to get the solution for 3x+ 3y= 10 -0x+ 9y= -30

OpenStudy (anonymous):

I don't get your equation. Is it like this: \[3x+3y=-30 and 10+9y=-30\]

OpenStudy (anonymous):

yes, sorry:)

OpenStudy (anonymous):

no wait, it's 9! sorry

OpenStudy (anonymous):

Ok, so you should multiply the equation with the threes to get a -9y so you can eliminate the y and solve for x

OpenStudy (anonymous):

by adding the equations. so -3( 3x+3y)=-3(-30)

OpenStudy (anonymous):

that doesn't make any sense, it would be a positive 9 and a negative nine both times, they'd equal zero

OpenStudy (anonymous):

I'm still not sure about the equations. Could you write them out again?

OpenStudy (anonymous):

sure. 3x+ 3y= 10 -9 -9y= -30

OpenStudy (anonymous):

perfect. Multiply the top equation thru by three to get\[9x+9y=30\]then add equations \[9x+9y=30\]\[-9-9y=-30\]to get \[9x-9=0\]\[9x=9\]\[x=1\]Substitute 1 for x in 9x+9y=30, and then solve for y. Y should equal 7/3, but check that in \[-9-9y=-30\]

OpenStudy (anonymous):

wow. thank you so much!

OpenStudy (anonymous):

yw